OpenStudy (anonymous):

does 200/15 = X/15 does X = 32 ?

OpenStudy (amistre64):

umm, no x would equal 200

OpenStudy (anonymous):

how ?

OpenStudy (amistre64):

multiply both sides by 15

OpenStudy (anonymous):

can u show mi what u mean by multipying both sides ?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

not really. Its pretty much as basic as i can get it

OpenStudy (amistre64):

id be curious to know why you think x=32 tho; where would you get that result from?

OpenStudy (anonymous):

cross mutliply and divied

OpenStudy (amistre64):

\[\frac{200}{15}=\frac{x}{15}\] \[\frac{200(15)}{15}=x\] like this?

OpenStudy (anonymous):

750/35 X/1000 does x = 1,000 ?

OpenStudy (amistre64):

or better yet: \[\frac{200}{15}=\frac{x}{15}\] \[200=\frac{x(15)}{15}\]

OpenStudy (amistre64):

your second problem makes no sense

OpenStudy (anonymous):

omg im soo sorry is dat im confuesed :/

OpenStudy (anonymous):

i need help in my 2nd problem